Suppose f(x)=2x-8 and g(x)=1x+11 find (f(-2) and g(f(-2))
Answer by tommyt3rd(5050)

f(-2)=2(-2)-8=4--8=-12

g(f(-2))=g(-12)= 1(-12)+11=-12+11=-1



:)